A 250 N force pushes a 25-kg box for 100.0 m. What work is done (by the push only) on the box?


If the box starts from rest and a 50 N frictional force resists the motion, what is the final speed of the box?


If the trup is up a 37 degree incline, what speed will it have after the 100 m?


1
Expert's answer
2021-02-15T17:44:27-0500

1) The work is defined as follows:


"W = Fs"

where "F = 250N" is the pushing force, and "s = 100m" is the displacement. Thus:


"W = 250N\cdot 100m = 25\times 10^{3}J = 25 \space kJ"

2) According to the energy-work theorem, the increase in the kinetic energy of the box is equal to the work of the net force acted on this box:


"K_2-K_1 = F_{net}s"

where "K_1 = 0" is the initial kinetic energy (since the box starts from rest), "F_{net} = 250N-50N = 200N" is the net foce acted on the box (pushing force minus resistance force), and


"K_2 = \dfrac{mv^2}{2}"

is the final kinetic energy, where "m = 25kg" is the mass, and "v" is the final speed of the box. Substituting and expressing "v", obtain:


"\dfrac{mv^2}{2}-0 = F_{net}s\\
v = \sqrt{\dfrac{2F_{net}s}{m}}\\
v = \sqrt{\dfrac{2\cdot 200\cdot 100}{25}} = 40m/s"

3) If the trup is up a "\theta =37\degree" incline and 100 m counts along it, the projection of the pushing force that actually pushes the box is:


"F = 250\cdot \cos37\degree \approx 199.66N"

Then the net force is:


"F_{net} = 199.66N-50N = 149.66N"

and the speed will be:


"v = \sqrt{\dfrac{2\cdot 149.66\cdot 100}{25}} \approx 34.6m/s"


Answer. 1) 25 kJ, 2) 40 m/s, 3) 34.6 m/s.